# Generative or Discriminative? Revisiting Text Classification in the Era of Transformers

### v1

@@ Abstract
The comparison between discriminative and generative classifiers has intrigued researchers since Efron (1975)’s seminal analysis of logistic regression versus discriminant analysis. While early theoretical work established that generative classifiers exhibit lower sample complexity but higher asymptotic error in simple linear settings, these trade-offs remain unexplored in the transformer era. We present the first comprehensive evaluation of modern generative and discriminative architectures—Auto-regressive, Masked Language Modeling, Discrete Diffusion, and Encoders for text classification. Our study reveals that the classical “two regimes" phenomenon manifests distinctly across different architectures and training paradigms. Beyond accuracy, we analyze sample efficiency, calibration, noise robustness, and ordinality across diverse scenarios. Our findings offer practical guidance for selecting the most suitable modeling approach based on real-world constraints such as latency and data limitations.
@@
Siva Rajesh Kasa, Karan Gupta, Sumegh Roychowdhury, Ashutosh Kumar, Yaswanth Biruduraju,
Santhosh Kumar Kasa, Nikhil Priyatam Pattisapu, Arindam Bhattacharya,
Shailendra Agarwal, Vijay huddar Amazon Inc. kasasiva,karaniis,sumegr@amazon.com
@@ 1 Introduction
Prior research on generative classifiers has largely focused on non-textual tabular data, utilizing linear models such as Linear Discriminant Analysis (Efron, 1975) and Naive Bayes (Ng and Jordan, 2001). While Yogatama et al. (2017) extended this analysis to neural architectures using RNNs and LSTMs (Hochreiter and Schmidhuber, 1997) for the TC task and found similar conclusions about generative advantages in low-data regimes, their study predated the transformer era. Modern NLP has seen the emergence of various successful transformer-based generative modeling paradigms, including auto-regressive (AR) models like GPT (Radford et al., 2018) that maximize likelihood directly, Discrete Diffusion models (Lou et al., 2024) that learn through iterative denoising, and masked language models (MLM) (Devlin et al., 2019) that optimize pseudo-likelihood Wang and Cho (2019). These approaches offer different trade-offs in terms of computational efficiency, sample complexity, and modeling flexibility. However, a systematic comparison of these paradigms for text classification remains unexplored, particularly in the context of varying model sizes and real-world deployment considerations.Our work fills this gap by providing a practitioner-oriented study that evaluates these approaches not just on classification accuracy, but also on crucial deployment metrics including different model scales, robustness to input perturbations, reliability of output probabilities through calibration analysis, and preservation of ordinal relationships between classes. This comprehensive evaluation aims to provide concrete guidance for choosing between different generative and discriminative approaches based on specific deployment constraints and requirements. We strategically focus on widely available public benchmark datasets for reproducibility purposes. Following Li et al. (2025) and Yogatama et al. (2017), our study evaluates all models trained from scratch, rather than relying on pre-trained weights, providing crucial insights for practitioners working with domain-specific data (Huang et al., 2019) or in resource-constrained environments (Martin et al., 2022). This approach helps isolate the confounding effects of the pre-training corpus (Razeghi et al., 2022) from other factors such as the modeling approach and size, which we evaluate. The combined effects can be explored in future work.
@@
Our main contributions include the following: (a) We present the first large-scale comparative study of two major classification approaches - Discriminative (Encoder) and Generative (Text Diffusion, AR, MLM) on 9 popular classification benchmark datasets, which is a first of its kind in the transformer era. Our study reveals a more nuanced interplay between model size and sample complexity than the previously known “two regimes” phenomenon. (b) We provide comprehensive analyses across multiple dimensions including model scaling behavior, sample efficiency, and performance in low-resource settings with models trained from scratch. We also introduce novel evaluation perspectives by examining ordinal relationships between classes, output calibration and robustness to input noise, offering insights beyond traditional classification metrics. (c) Finally, we provide practical recommendations in Section 6 on selecting the appropriate model for deployment in various real-world scenarios.
@@ 2 Related Work
Pseudo-Generative Models. Recent work Sahoo et al. (2024) highlights a natural connection between Discrete Text Diffusion Lou et al. (2024) and the Masked Language Modeling (MLM) objective in BERT Devlin et al. (2019), showing that the diffusion objective can be expressed as a weighted sum of MLM losses. Using transformer encoder models, this approach achieves likelihood bounds comparable to or better than those in Lou et al. (2024). Motivated by this, we include vanilla MLM as a baseline for text classification. While MLM has typically served as a pretraining objective followed by fine-tuning Liu et al. (2019), there has been little systematic study of its direct use for classification. Although MLM does not explicitly model $P(X|y)$, it estimates $P(x_{m}|x_{\setminus m})$, where $x_{m}$ is a masked token and $x_{\setminus m}$ represents all other tokens. This approximates the pseudo-likelihood of $P(X,y)$ when modeled over the corpus (Wang and Cho, 2019). We therefore classify MLM as a pseudo-generative model.
Also, traditional generative classifiers aim to model $P(X|y)$ by prepending the label token. However, recent work Li et al. (2025) shows that appending the label at the end—though not strictly modeling $P(X|y)$—can yield better in-distribution performance. This setup also enables efficient inference, requiring only a single forward pass to predict the label, unlike traditional generative models that need #$\#_{\text{label}}$ forward passes. These benefits motivate the inclusion of such pseudo-generative models in our benchmarks. Notably, these approaches involve minimal changes to standard transformer architectures—typically just altering label placement or the loss function—while preserving the core model design. This allows for fair comparisons using widely available implementations accessible to practitioners.
@@ 3 Methodology
We approach the problem of label classification by leveraging two popular language modeling paradigms: (a) Generative - Discrete Diffusion models, Auto-regressive models (AR), and Masked Language Models (MLM) & (b) Discriminative - Encoder-based transformer models. Note that, for brevity, we use the term “generative” from this point onward to also include the pseudo-generative baselines. Let $\mathcal{D}=\{(X_{i},y_{i})\}_{i=1}^{N}$ denote the dataset where $X_{i}$ is the input text and $y_{i}\in\mathcal{Y}$ is the corresponding label from a finite set of classes $\mathcal{Y}$. Generative models tend to learn the joint data distribution $P(X,y)$ first and then try to infer the label using the marginals, whereas Discriminative models directly learn the conditional distribution $P(y|X)$. Note that each $X_{i}=x_{i}^{1}\ldots x_{i}^{n}$, where $x_{i}^{j}$ is a token from the associated vocabulary $\mathcal{V}$.
@@ 3.1 Discriminative Model for Classification
$\hat{y}_{i}=\texttt{softmax}(Wh_{i}),\quad\mathcal{L}_{\text{enc}}=-\sum_{i=1}^{N}\log P(y_{i}|X_{i})$ (1)
@@ 3.2 Generative Models for Classification
(2) Masked Language Modeling (MLM): During training, we mask 15% of tokens in input sequence $X_{i}=x_{i}^{1}\ldots x_{i}^{n}$ following Devlin et al. (2019) and predict them using unmasked bi-directional context. Wang and Cho (2019) show that the MLM objective stochastically captures the pseudo-loglikelihood which makes it similar to a denoising autoencoder Vincent et al. (2010). Hence, we consider MLM under the generative family of models. Formally, the objective is:
$\mathcal{L}_{\text{mlm}}=-\sum_{i=1}^{N}\sum_{j\in\mathcal{M}_{i}}\log P(x_{i}^{j}|X_{i}^{\setminus j})$ (2)
where $\mathcal{M}_{i}$ is the set of masked positions and $X_{i}^{\setminus j}$ denotes the unmasked input with only token at position $j$ masked. At inference, we use the template:
[…]
and predict the masked label token. The output vocabulary is restricted to the label token set $\mathcal{V}_{\mathcal{Y}}$.
@@
$\quad-\frac{p(z\mid x_{0})}{p(x\mid x_{0})}\log s_{\theta}(x)_{z}\Big{)}\bigg{]}$ (6)
@@
$-\log p^{\theta}_{0}(x_{0})\leq\mathcal{L}_{DWDSE}(x_{0})+constant$ (7)
@@ 4.2 Experimental Setup
[…]
In the second part of our evaluation, we assess each model’s robustness to input perturbations. In real-world scenarios, particularly in e-commerce platforms, often encounter various text corruptions (like OCR errors in product documentation, truncated reviews, or incomplete user queries), we focus on two systematic types of synthetic noise to evaluate model robustness: (a) Random Token Drop — where Xof tokens are randomly removed from the input sentence, and (b) Random Token Substitution — where Xof tokens are replaced with random tokens from the vocabulary (excluding special tokens like [PAD], [MASK]). We conduct these experiments to explore Q2 about model robustness to input perturbations.
@@ 5 Results
Config ENC ARpseudo AR MLM DIFF
(12L,12H) 51.1% 46.7% 37.8% 34.4% 54.4%
(6L,6H) 47.8% 51.1% 46.7% 46.7% 53.3%
Table 1: Minimum noisefor 10% $\Delta$ weighted-F1 drop under Random Token Dropping. ($\uparrow$ is better)
Config ENC ARpseudo AR MLM DIFF
(12L,12H) 32.2% 22.2% 28.9% 31.1% 35.6%
(6L,6H) 37.8% 21.1% 30.0% 32.2% 34.4%
Table 2: Minimum noisefor 10% $\Delta$ weighted-F1 drop under Random Token Substitution. ($\uparrow$ is better)
@@
Figure 3 shows that ARpseudo generally underperforms AR and also displays higher variance in low-data settings—the recommended use case—while the opposite holds in high-data scenarios. This reveals a new insight beyond Li et al. (2025), who only evaluated full-data settings where ARpseudo performed better in-distribution. As noted in Section 3, AR requires $|label|$-times forward passes per prediction, unlike the single pass needed for ARpseudo; however, this can be mitigated via batching or parallel processing, reducing inference time differences at the cost of higher computation.
@@
Q2: We evaluate the robustness of all approaches under both 6-layer and 12-layer configurations across two noise schedules in full-data settings. We exclude 1-layer models from this analysis since their performance is mostly trivial (except for ENC), making robustness comparisons uninformative. In Tables 1 and 2, we report the minimum noise level at which each model’s performance drops by 10% $\Delta$ (similar trend for 5% & 15%) relative to its peak, averaged across all datasets, as a measure of robustness boundary. Our analysis reveals that all models exhibit lower robustness to substitution noise compared to dropping. This can be explained by the inequality: $P(\text{garbage}_{t}|X_{1...t-1})<P(x_{t+1}|X_{1...t-1})$—the model is more likely to assign lower probability to a corrupted token than to a skip token at $t+1$-th position (assuming $t$-th token was dropped), which may still be contextually relevant given $X_{1...t-1}$.
@@
The generative DIFF demonstrates superior robustness to both token dropping and substitution (except in 6-layers where ENC is slightly better), likely because its training paradigm involves recovering true tokens from noise/masked inputs. The discriminative ENC maintains consistent robustness under both noise types, while generative AR shows the high sensitivity to noise. Combining these findings with Q1’s results reveals that generative AR models face dual challenges compared to ENC in full data settings: they underperform in terms of both weighted-f1 and robustness. This contrasts with Li et al. (2025)’s findings that discriminative ENC models rely on shortcuts and show less robustness compared to generative AR. However, their analysis focuses on shortcut learning and distribution shifts rather than input perturbation noise across varying model sizes. Notably, while the pseudo-generative MLM and ARpseudo demonstrates superior performance in larger models at full data settings, they exhibit lower robustness compared to similarly performing ENC models. Moreover the relative drop in robustness in moving from dropping to substitution noise is more severe in ARpseudo compared to AR. This is likely because ARpseudo conditions on corrupted inputs, so it’s directly affected by garbage tokens polluting the predictive context. However, for AR clean label conditions the model, and the noisy input is scored globally — giving the model more flexibility to discount garbage.

### v2

@@ Abstract
The comparison between discriminative and generative classifiers has intrigued researchers since Efron (1975)’s seminal analysis of logistic regression versus discriminant analysis. While early theoretical work established that generative classifiers exhibit lower sample complexity but higher asymptotic error in simple linear settings, these trade-offs remain unexplored in the transformer era. We present the first comprehensive evaluation of modern generative and discriminative architectures—Auto-regressive, Masked Language Modeling, Discrete Diffusion, and Encoders for text classification. Our study reveals that the classical “two regimes" phenomenon manifests distinctly across different architectures and training paradigms. Beyond accuracy, we analyze sample efficiency, calibration, noise robustness, and ordinality across diverse scenarios. Our findings offer practical guidance for selecting the most suitable modeling approach based on real-world constraints such as latency and data limitations.  [^1] Code available at: https://github.com/amazon-science/Generative-vs-Discriminative-Classifiers
@@
Siva Rajesh Kasa, Karan Gupta, Sumegh Roychowdhury, Ashutosh Kumar,
Yaswanth Biruduraju, Santhosh Kumar Kasa, Nikhil Priyatam Pattisapu,
Arindam Bhattacharya, Shailendra Agarwal, Vijay Huddar Amazon Inc. kasasiva,karaniis,sumegr@amazon.com
@@ 1 Introduction
Prior research on generative classifiers has largely focused on non-textual tabular data, utilizing linear models such as Linear Discriminant Analysis (Efron, 1975) and Naive Bayes (Ng and Jordan, 2001). While Yogatama et al. (2017) extended this analysis to neural architectures using RNNs and LSTMs (Hochreiter and Schmidhuber, 1997) for the TC task and found similar conclusions about generative advantages in low-data regimes, their study predated the transformer era. Modern NLP has seen the emergence of various successful transformer-based generative modeling paradigms, including auto-regressive (AR) models like GPT (Radford et al., 2018) that maximize likelihood directly, Discrete Diffusion models (Lou et al., 2024) that learn through iterative denoising, and masked language models (MLM) (Devlin et al., 2019) that optimize pseudo-likelihood Wang and Cho (2019). These approaches offer different trade-offs in terms of computational efficiency, sample complexity, and modeling flexibility. However, a systematic comparison of these paradigms for text classification remains unexplored, particularly in the context of varying model sizes and real-world deployment considerations. Our work fills this gap by providing a practitioner-oriented study that evaluates these approaches not just on classification accuracy, but also on crucial deployment metrics including different model scales, robustness to input perturbations, reliability of output probabilities through calibration analysis, and preservation of ordinal relationships between classes. This comprehensive evaluation aims to provide concrete guidance for choosing between different generative and discriminative approaches based on specific deployment constraints and requirements. We strategically focus on widely available public benchmark datasets for reproducibility purposes. Following Li et al. (2025) and Yogatama et al. (2017), our study evaluates all models trained from scratch, rather than relying on pre-trained weights, providing crucial insights for practitioners working with domain-specific data (Huang et al., 2019) or in resource-constrained environments (Martin et al., 2022). This approach helps isolate the confounding effects of the pre-training corpus (Razeghi et al., 2022) from other factors such as the modeling approach and size, which we evaluate.
@@
Our main contributions include the following: (a) We present the first large-scale comparative study of two major classification approaches - Discriminative (Encoder) and Generative (Text Diffusion, AR, MLM) on 9 popular classification benchmark datasets, which is a first of its kind in the transformer era. Our study reveals a more nuanced interplay between model size and sample complexity than the previously known “two regimes” phenomenon. (b) We provide comprehensive analyses across multiple dimensions including model scaling behavior, sample efficiency, and performance in low-resource settings with models trained from scratch. We also introduce novel evaluation perspectives by examining ordinal relationships between classes, output calibration and robustness to input noise, offering insights beyond traditional classification metrics. We also evaluate these paradigms using pretrained models. (c) Finally, we provide practical recommendations in Section 6 on selecting the appropriate model for deployment in various real-world scenarios, a concise summary of which is given in Table 1.
@@
Properties ENC AR AR(p) MLM DIFF
[…]
Table 1: Comparison of different classification approaches across key properties. ENC: Encoder-based classification, AR: Auto-Regressive Model, AR(p): Pseudo-AR model, MLM: Masked Language Modeling, DIFF: Discrete Text Diffusion. Values indicate relative performance/requirements (High/Medium/Low). $\blacksquare$ indicates preferred characteristics, $\blacksquare$ indicates less favorable ones.
@@ 2 Related Work
Pseudo-Generative Models. Recent work Sahoo et al. (2024) highlights a natural connection between Discrete Text Diffusion Lou et al. (2024) and the Masked Language Modeling (MLM) objective in BERT Devlin et al. (2019), showing that the diffusion objective can be expressed as a weighted sum of MLM losses. Using transformer encoder models, this approach achieves likelihood bounds comparable to or better than those in Lou et al. (2024). Motivated by this, we include vanilla MLM as a baseline for text classification by first appending ‘‘The label is y" as a suffix during training and appending ‘‘The label is [MASK]" as a suffix during inference. While MLM has typically served as a pretraining objective followed by fine-tuning Liu et al. (2019), there has been little systematic study of its direct use for classification. Although MLM does not explicitly model $P(X|y)$, it estimates $P(x_{m}|x_{\setminus m})$, where $x_{m}$ is a masked token and $x_{\setminus m}$ represents all other tokens. This approximates the pseudo-likelihood of $P(X,y)$ when modeled over the corpus (Wang and Cho, 2019). We therefore classify MLM as a pseudo-generative model.
Also, traditional generative classifiers aim to model $P(X|y)$ by prepending the label token. However, recent work Li et al. (2025) shows that appending the label at the end—though not strictly modeling P(X|y)P(X|y)—can yield better in-distribution performance. This setup also enables efficient inference, requiring only a single forward pass to predict the label, unlike traditional generative models that need $_{\text{label}}$ forward passes. These benefits motivate the inclusion of such pseudo-generative models in our benchmarks. Notably, these approaches involve minimal changes to standard transformer architectures—typically just altering label placement or the loss function—while preserving the core model design. This allows for fair comparisons using widely available implementations accessible to practitioners.
@@ 3 Methodology
We approach the problem of TC by leveraging two popular language modeling paradigms: (a) Generative - Discrete Diffusion models, Auto-regressive models (AR), and Masked Language Models (MLM) & (b) Discriminative - Encoder-based transformer models. Note that, for brevity, we use the term “generative” from this point onward to also include the pseudo-generative baselines. Let $\mathcal{D}=\{(X_{i},y_{i})\}_{i=1}^{N}$ denote the dataset where $X_{i}$ is the input text and $y_{i}\in\mathcal{Y}$ is the corresponding label from a finite set of classes $\mathcal{Y}$. Generative models tend to learn the joint data distribution $P(X,y)$ first and then try to infer the label using the marginals, whereas Discriminative models directly learn the conditional distribution $P(y|X)$. Note that each $X_{i}=x_{i}^{1}\ldots x_{i}^{n}$, where $x_{i}^{j}$ is a token from the associated vocabulary $\mathcal{V}$.
@@ 3.1 Discriminative Model for Classification
$\hat{y}_{i}=\texttt{softmax}(Wh_{i}),\mathcal{L}_{\text{enc}}=-\sum_{i=1}^{N}\log P(y_{i}|X_{i})$
@@ 3.2 Generative Models for Classification
(2) Masked Language Modeling (MLM): During training, we first modify the input $X_{i}$ to :
$X^{\prime}_{i}=\texttt{[CLS]}\;X_{i}\;\texttt{[SEP]}\;\texttt{``The label is"}\;$
We then apply the standard mask i.e., on 15% of tokens in input sequence $X^{\prime}_{i}\oplus y_{i}=x_{i}^{1}\ldots x_{i}^{n^{\prime}}y_{i}$ following Devlin et al. (2019) and predict them using unmasked bi-directional context. Wang and Cho (2019) show that the MLM objective stochastically captures the pseudo-loglikelihood which makes it similar to a denoising autoencoder Vincent et al. (2010). Hence, we consider MLM under the generative family of models. Formally, the objective is:
$\mathcal{L}_{\text{mlm}}=-\sum_{i=1}^{N}\sum_{j\in\mathcal{M}_{i}}\log P(x_{i}^{j}|X^{\prime}_{i}\oplus y_{i}^{\setminus j})$ (1)
where $\mathcal{M}_{i}$ is the set of masked positions and $X^{\prime}_{i}\oplus y_{i}^{\setminus j}$ denotes the unmasked input with only token at position $j$ masked. At inference, we use the template:
$X^{\prime}_{i}=\texttt{[CLS]}\;X_{i}\;\texttt{[SEP]}\;\texttt{"The label is"}\;\texttt{[MASK]}\;\texttt{.}$
and predict the masked label token. The output vocabulary is restricted to the label token set $\mathcal{V}_{\mathcal{Y}}$. Since MLM returns token probabilities across the entire vocabulary for a [MASK] token, we extract the dimensions corresponding to the label tokens and normalize them to sum to 1, thereby obtaining the class probabilities.
@@
$-\frac{p(z\mid x_{0})}{p(x\mid x_{0})}\log s_{\theta}(x)_{z}\Big)\bigg]$ (5)
@@
$\log p^{\theta}_{0}(x_{0})\leq\mathcal{L}_{DWDSE}(x_{0})+constant$ (6)
@@ 4.2 Experimental Setup
We experiment with multiple dataset sample sizes $\in$ {128, 256, 512, 1024, 2048, 4096, full_data}. To assess the effect of model sizes, we test 3 model size configurations using the base Transformer architecture: small (1 layer, 1 head), medium (6 layers, 6 heads) and large (12 layers, 12 heads). Performance is measured using the weighted-F1 score. All experiments are repeated with 3 random seeds, running a total of $9\times 7\times 3\times 3\times 5=2835$ experiments and we report the average and shaded standard deviations in Figure 23. These experiments are designed to address Q1
In the second part of our evaluation, we assess each model’s robustness to input perturbations. In real-world scenarios, particularly in e-commerce platforms, often encounter various text corruptions (like OCR errors in product documentation, truncated reviews, or incomplete user queries), we focus on two systematic types of synthetic noise to evaluate model robustness: (a) Random Token Drop — where X% of tokens are randomly removed from the input sentence, and (b) Random Token Substitution — where X% of tokens are replaced with random tokens from the vocabulary (excluding special tokens like [PAD], [MASK]). We conduct these experiments to explore Q2 about model robustness to input perturbations.
@@ 5 Results
Config Metric 5% 10% 15% 20% 30%
[…]
AR-pseudo 33.3 46.7 61.1 73.3 86.7
AR 25.6 37.8 50.0 67.8 86.7
MLM 23.3 34.4 47.8 61.1 71.1
[…]
Table 2: Minimum noise% needed for X% weighted-F1 drop from the peak under Random Token Dropping. ($\uparrow$ is better)
Config Metric 5% 10% 15% 20% 30%
[…]
AR-pseudo 15.6 21.1 27.8 32.2 50.0
AR 21.1 30.0 38.9 47.8 62.2
[…]
AR-pseudo 13.3 22.2 27.8 34.4 51.1
[…]
Table 3: Minimum noise% needed for X% weighted-F1 drop from the peak under Random Token Substitution. ($\uparrow$ is better)
@@
Figure 3 shows that ARpseudo generally underperforms AR and also displays higher variance in low-data settings—the recommended use case—while the opposite holds in high-data scenarios. This reveals a new insight beyond Li et al. (2025), who only evaluated full-data settings where ARpseudo performed better in-distribution. As noted in Section 3, AR requires $|label|$-times forward passes per prediction, unlike the single pass needed for ARpseudo; however, this can be mitigated via batching or parallel processing, reducing inference time differences at the cost of higher computation. We also investigate the claim that “larger models can sometimes deteriorate performance” (Nakkiran et al., 2019) in the Appendix G. While we observe the classical bias-variance trade-off in small-data regimes, performance generally improves with model size in full-data settings especially for AR.
@@
Q2: We evaluate the robustness of all approaches under both 6-layer and 12-layer configurations across two noise schedules in full-data settings. We exclude 1-layer models from this analysis since their performance is mostly trivial (except for ENC), making robustness comparisons uninformative. In Tables  2 and  3, we report the minimum noise level required to degrade a model’s performance by a certain threshold $X\%=\{5\%,10\%,\dots\}$ relative to its peak, averaged across all datasets, as a measure of robustness boundary. Our analysis reveals that all models exhibit lower robustness to substitution noise compared to dropping. This can be explained by the inequality: $P(\text{garbage}_{t}|X_{1...t-1})<P(x_{t+1}|X_{1...t-1})$—the model is more likely to assign lower probability to a corrupted token than to a skip token at $t+1$-th position (assuming $t$-th token was dropped), which may still be contextually relevant given $X_{1...t-1}$.
@@
The generative DIFF demonstrates superior robustness to both token dropping and substitution (except in 6-layers where ENC is slightly better), likely because its training paradigm involves recovering true tokens from noise/masked inputs. The discriminative ENC maintains consistent robustness under both noise types, while generative AR shows the high sensitivity to noise. Combining these findings with Q1’s results reveals that generative AR models face dual challenges compared to ENC in full data settings: they underperform in terms of both weighted-f1 and robustness. This contrasts with Li et al. (2025)’s findings that discriminative ENC models rely on shortcuts and show less robustness compared to generative AR. However, their analysis focuses on shortcut learning and distribution shifts rather than input perturbation noise across varying model sizes. Notably, while the pseudo-generative MLM and ARpseudo demonstrate superior performance in larger models at full data settings, they exhibit lower robustness compared to similarly performing ENC models. Moreover the relative drop in robustness in moving from dropping to substitution noise is more severe in ARpseudo compared to AR. This is likely because ARpseudo conditions on corrupted inputs, so it’s directly affected by garbage tokens polluting the predictive context. However, for AR clean label conditions the model, and the noisy input is scored globally — giving the model more flexibility to discount garbage.
@@
5.1 Impact of Initialization with Pretrained models
To investigate the impact of pretraining, we conducted additional experiments using BERT-base (for ENC) and GPT-2 base (for AR), both featuring comparable architectures with 12 layers and 12 attention heads. These models were fine-tuned on our benchmark datasets across various sample sizes, maintaining consistency with our previous experimental protocol. Figure 10 has the detailed results.
These experiments reveal important insights that contrast with our findings from models trained from scratch. When using pretrained weights, we observe that the classical “two regimes" phenomenon no longer holds. Instead, the discriminative ENC model consistently outperforms the generative AR approach across all data regimes in most datasets. This aligns with recent findings from Zheng et al. (2023) in the vision domain, where pretraining was shown to eliminate the two-regime effect. This behavior can be theoretically explained by viewing pretraining as providing models with “asymptotically large" amounts of data, effectively reducing the traditional advantage that generative models hold in low-data settings since both architectures begin with rich, generalized representations. However, these results should be interpreted with several caveats in mind: pretrained models often employ mixed training objectives (e.g., BERT uses both MLM and Next Sentence Prediction (NSP)), rely on different pretraining datasets with varying cutoff dates, and have distinct architectural designs. Additionally, our pretrained analysis was limited to comparing AR and ENC models due to the current unavailability of pretrained diffusion models.